# Fluttering without wind: Stokesian quasiperiodic settling

### v1

@@ Abstract
Another class of periodic orbits are obtained on the center fixed points of the $\bm{p}_{y}$ dynamics, characterized by $H=H_{\text{max}}$. In this case $\bm{A}(t)=\bm{A}^{*}$ is a constant matrix with the eigenvalues $0,\pm i\sqrt{\alpha_{p}\alpha_{r}}$. The eigenvector corresponding to the zero eigenvalue is not orthogonal to $\bm{p}_{y}$, and is dropped from the discussion since it violates the constraints.
The translational velocity of any of these three classes of bodies can be obtained from equations (1) and (5). In the lab frame
$v_{x}=-G_{1}p_{1x}p_{1y}-G_{2}p_{2x}p_{2y}-G_{3}p_{3x}p_{3y},$ (15a)
$v_{y}=-G_{1}p_{1y}^{2}-G_{2}p_{2y}^{2}-G_{3}p_{3y}^{2},$ (15b)
$v_{z}=-G_{1}p_{1z}p_{1y}-G_{2}p_{2z}p_{2y}-G_{3}p_{3z}p_{3y},$ (15c)
$G_{1}\equiv\frac{C_{3}}{A_{1}C_{3}-B_{31}^{2}},\quad G_{3}\equiv\frac{C_{1}}{A_{3}C_{1}-B_{13}^{2}}.$ (16)
Viscous dissipation requires $G_{1},G_{2},G_{3}>0$ [4, 5]. Equation (Fluttering without wind: Stokesian quasiperiodic settling) shows direct dependence of translational velocity on the orientation of the body. Thus, periodicity in ${\mathbf{p}}_{y}$ implies periodicity in $v_{y}$, and quasi-periodicity in ${\mathbf{p}}_{x}$ implies quasi-periodicity in the horizontal components of the velocity. It is relevant to ask if there a persistent drift in the horizontal plane. This must be answered separately for the quasi-periodic and the periodic cases. For the former, following Thorp and Lister [14] we exploit the discrete symmetries of the rotational dynamical equations to rule out persistent drift. Note that $v_{x}(\bm{p}_{y},\bm{p}_{x})=-v_{x}(\bm{p}_{y},-\bm{p}_{x})$, $v_{z}(\bm{p}_{y},\bm{p}_{x})=-v_{z}(\bm{p}_{y},-\bm{p}_{x})$, a quasi-periodic trajectory which explores $(\bm{p}_{y},\bm{p}_{x})$ must explore $(\bm{p}_{y},-\bm{p}_{x})$ (see appendix C). Thus, any drift at one point in time is exactly cancelled by the opposite drift at some other point in time. For the periodic case with $H\neq H_{\text{max}}$, persistent drift can occur in the horizontal plane only if the response frequency $\mu$ is exactly equal to the driving frequency $\omega_{y}$ (see appendix C). In other words, the Poincare map $\mathcal{P}_{t_{0}}^{t_{0}+T_{y}}$ now is the identity matrix. When $H=H_{\text{max}}$, we will still have no persistent drift, since the valid solutions of $\mathbf{p}_{x}$ are periodic with zero mean. Finally, for $H=0$, the body exponentially reaches the saddle point $\bm{p}_{y}^{*}=(0,\pm 1,0)$ along its stable manifold as $t\to\infty$ and consequently the horizontal drift goes to zero exponentially.
Figure 7: Translational trajectories in the horizontal plane for sample (a) quasi-periodic and (b) periodic orbits. With gravity pointing out of the plane, the trajectories move clockwise in this view, which we classify as negative chirality.
We now discuss the chirality of the translational trajectories. As the body settles, its motion in the horizontal $x-z$ plane is either clockwise or anti-clockwise when viewed from below, corresponding to negative or positive chirality respectively. As mentioned in figure 3, regions on $\mathcal{S}^{2}$ with $p_{1y}p_{3y}>0$ correspond to different chirality in translation motion from regions with $p_{1y}p_{3y}<0$. This can be seen as follows: for any given angular trajectory ($\bm{p}_{y},\bm{p}_{x}$) with $p_{1y}p_{3y}>0$, the transformation $g:(p_{1y},\,p_{1x})\to(-p_{1y},\,-p_{1x})$ maps the original trajectory to the trajectory in the region $p_{1y}p_{3y}<0$ with the same $H$. Thus the mapped trajectory has same time period $T_{y}$ in $\bm{p}_{y}$ dynamics and same frequency ratio $\mu/\omega_{y}$ in $\bm{p}_{x}$ dynamics. However, $g:v_{z}\to-v_{z}$, while the other velocity components remain invariant. So the trajectory of the translational dynamics ($x(t),y(t),z(t)$) is mapped to the trajectory $(x(t),y(t),-z(t))$, i.e., the original and mapped trajectories are of opposite chirality.
The vertical sedimentation velocity $v_{y}$ is $T_{y}$-periodic because it only depends on $\bm{p}_{y}$. Thus, for a given body the mean sedimentation velocity $\langle v_{y}\rangle$ averaged over $T_{y}$ only depends on $H$. For $H=0$, $v_{y}$ reaches $v_{y0}=-G_{2}$ exponentially fast in time with the rate $|\alpha_{p}|$ or $|\alpha_{r}|$, depending on the initial condition, and for $H=H_{\text{max}}$, $v_{y}=v_{ym}\equiv-(G_{1}\alpha_{p}+G_{3}\alpha_{r})/(\alpha_{p}+\alpha_{r})$. The shape of the body determines whether $\langle v_{y}\rangle$ increases or decreases with $H$, and a sample of each is shown in figure 8.
Figure 8: Dependence of mean vertical sedimentation velocity on $H$. Depending on the shape of the body, the periodic/quasi-periodic dynamics $H\neq 0$ can result in either slower or faster settling velocity as compare to the case where it falls along a stable manifold of saddle points $H=0$.
@@
It remains to obtain the resistance matrix $\cal R$, and thence $\alpha_{p}$ and $\alpha_{r}$, for a given body shape. To define the set of bodies we have used two shape-controlling non-dimensional parameters: the bending parameter $a$, and the concavity parameter $b$, as shown in figure 9. Further details are available in supplementary E.
Figure 9: Our set of bodies is defined by the two parameters $a$ and $b$ shown here. The body shown is a flutterer.
We develop and use a numerical solver to find the resistance matrix $\cal R$ by the Boundary Integral Method [5, 17, 18, 19, 20]. The five parameters $\alpha_{p},\,\alpha_{r},\,G_{1},\,G_{2},\,G_{3}$ completely describe the sedimentation problem in a quiescent fluid.
@@
Figure 10: As $a$ and $b$ in figure 9 are varied, we obtain settlers ($\alpha_{p}\alpha_{r}<0$), drifters ($\alpha_{p}\alpha_{r}=0$), and flutterers ($\alpha_{p}\alpha_{r}>0$). (a) Increasing $b$ turns settlers into flutterers, but increasing $a$ at fixed $b$ eventually turns flutterers back into settlers. (b) Dependence of parameters determining the settling velocity on the bending parameter $a$, for a fixed $b=0.4$.
To summarise, we have shown that bodies with exactly two planes of symmetry show three kinds of Stokesian sedimentation, depending only on their shape, appearing through the parameter $\alpha_{p}\alpha_{r}$. (1) Settlers with $\alpha_{p}\alpha_{r}<0$ asymptotically fall vertically without rotating, (2) Drifters with $\alpha_{p}\alpha_{r}=0$ eventually fall obliquely with respect to gravity without rotating and (3) Flutterers with $\alpha_{p}\alpha_{r}>0$ perform periodic or quasi-periodic rotational and translational motion as they sediment. The vertical projections $\bm{p}_{y}$ of their body-fixed coordinate axes executes periodic dynamics. A conserved quantity $H$ exists, the value of which decides the time period of the $\bm{p}_{y}$ dynamics. The periodic motion in $\bm{p}_{y}$ drives the dynamics of horizontal projections of body attached coordinate axes $\bm{p}_{x}$, which leads to a periodic or quasi-periodic response based on the ratio of response to driving frequency, $\mu/\omega_{y}$. This ratio too depends only on $H$. Our numerical solutions suggest that this ratio is bounded within the interval $[1-1/\sqrt{2},0.5]$, indicating that no value of $H$ corresponds to the response frequency being same as the driving frequency. For these periodic or quasi-periodic orbits there is no persistent drift in the horizontal plane. The chirality of the translational trajectories can be classified using $p_{1y}p_{3y}\lessgtr 0$. The mean sedimentation velocity of the flutterers has the same time period as the $\bm{p}_{y}$ dynamics. All three classes of bodies can be obtained by bending a capsule about its central axis and making its cross-section concavo-convex. This work highlights how rich the dynamics of sedimenting bodies can be in the presence of even simple translation-rotation coupling. Our choice of dynamical variables brings out an inherent decoupling in the rotational dynamics, providing better understanding of flutterers, and can be used to study them in background flow.
@@
We thank Saumav Kapoor for help in designing the flutterer shape. We acknowledge support of the Department of Atomic Energy, Government of India, under project no. RTI4001.
@@ Appendix B Discrete and continuous symmetries of the rotational dynamics of flutterers
The discrete symmetry group of the rotational dynamical equations (Fluttering without wind: Stokesian quasiperiodic settling) and (Fluttering without wind: Stokesian quasiperiodic settling) is $\mathcal{G}_{R}=\mathbb{Z}_{2}\times\mathbb{Z}_{2}\times\mathbb{Z}_{2}$. The group $\mathcal{G}_{R}$ is an abelian group containing 8 elements which corresponds to the following eight transformations of dynamical variables:
@@
The conserved quantity $H$ is obtained as a first integral using: $\bm{Q}_{1}\cdot\bm{\Lambda}_{2}=\alpha_{p}+\alpha_{r}+t\dfrac{d}{dt}\log(p_{1y}^{\alpha_{p}}p_{3y}^{\alpha_{r}})$. The other scalar products of $\bm{Q}$’s and $\bm{\Lambda}$’s yield the constraints (Fluttering without wind: Stokesian quasiperiodic settling) as the conserved quantities.
Appendix C Necessary condition for persistent horizontal drift
@@
Claim: Flutterers can have persistent drift in the plane perpendicular to the gravity only if the driving frequency $\omega_{y}$ of the $\bm{p}_{y}$ dynamics is exactly equal to the response frequency $\mu$ of the $\bm{p}_{x}$ dynamics, i.e. $\mu=\omega_{y}$. In this case, the Poincare map $\mathcal{P}_{t_{0}}^{t_{0}+T_{y}}$ is just the identity map.
@@
Proof: The rotational dynamical equations (Fluttering without wind: Stokesian quasiperiodic settling) and (Fluttering without wind: Stokesian quasiperiodic settling) have a discrete symmetry group $\mathcal{G}_{R}=\mathbb{Z}_{2}\times\mathbb{Z}_{2}\times\mathbb{Z}_{2}$. If a given rotational trajectory is invariant under $g\in\mathcal{G}_{R}$ and $g(v_{i})=-v_{i}$ then there can be no persistent drift along the $x_{i}$ direction for that trajectory, for the drift at one point is exactly cancelled by the other mapped point. Note that $g:(p_{1y},p_{2y},p_{3y},p_{1x},p_{2x},p_{3x})\to(p_{1y},p_{2y},p_{3y},-p_{1x},-p_{2x},-p_{3x})$ is a symmetry of equations (Fluttering without wind: Stokesian quasiperiodic settling) and (Fluttering without wind: Stokesian quasiperiodic settling) with $g(v_{x})=-v_{x}$ and $g(v_{z})=-v_{z}$. The invariance of trajectories of the rotational dynamical system under $g$ requires the inversion symmetry $\bm{p}_{x}\to-\bm{p}_{x}$.
@@
For quasi-periodic cases, any point $\bm{p}_{x}(t)$ in a given trajectory will be mapped to the point $-\bm{p}_{x}$ at time $t+kT_{y},\,k\in\mathbb{N}$. This is due to the fact that the Poincare map $\mathcal{P}_{t}^{t+T_{y}}$ fills up the entire great circle with normal vector $\bm{p}_{y}(t)$. Since this great circle has an inversion symmetry, there exist a $k\in\mathbb{N}$ such that $\bm{p}_{x}(t+kT_{y})=-\bm{p}_{x}$. Therefore, the region occupied by quasi-periodic orbits on $S^{2}$ has inversion symmetry $\bm{p}_{x}\to-\bm{p}_{x}$. This rules out any persistent drift in the horizontal plane owing to the discrete symmetry element $g\in\mathcal{G}_{R}$.
@@
For the periodic case we have closed trajectories subject to constraints (Fluttering without wind: Stokesian quasiperiodic settling) which may not have the inversion symmetry $\bm{p}_{x}\to-\bm{p}_{x}$ (see figure 6). The discrete symmetry argument to rule out persistent drift is inconclusive in this case. A necessary condition for persistent drift can be obtained as follows: Owing to the equation (14), each term in the right hand side of $v_{x}$ and $v_{z}$ in equation (Fluttering without wind: Stokesian quasiperiodic settling) is of the form $e^{i\mu t}f(t)$, where $f(t)=f(t+T_{y})$ is some $T_{y}$ periodic function. The contribution of such terms to the displacement over one time period $T$ of a periodic trajectory is
@@
Numerics suggests that the ratio $\mu/\omega_{y}$ is bounded between the interval [11/2, $[1-1/\sqrt{2},\,0.5]$ in which case there is no persistent drift.

### v2

@@ Abstract
Another class of periodic orbits are obtained on the centre fixed points of the $\bm{p}_{y}$ dynamics, characterized by $H=H_{\text{max}}$. In this case $\bm{A}(t)=\bm{A}^{*}$ is a constant matrix with the eigenvalues $0,\pm i\sqrt{\alpha_{p}\alpha_{r}}$. The eigenvector corresponding to the zero eigenvalue is not orthogonal to $\bm{p}_{y}$, and is dropped from the discussion since it violates the constraints. In this case, the body simply rotates about gravity and falls in a helical trajectory [6].
The translational velocity of di-bilaterals in the lab frame depends directly on $\bm{p}_{x}$ and $\bm{p}_{y}$ (see Supplemental Material). Thus, the translational trajectories inherit the periodicity or quasi-periodicity from the angular dynamics. A flutterer does not display persistent horizontal drift and maintains the same sign of chirality (see figure 7) throughout (more details in the Supplemental Material).
Figure 7: Sample trajectories in the horizontal plane for (a) quasi-periodic and (b) periodic orbits. The black arrows indicate the chirality of the settling trajectories.
@@
It remains to obtain the resistance matrix $\cal R$, and thence $\alpha_{p}$ and $\alpha_{r}$, for a given di-bilateral. We design a set of bodies with two parameters: the bending parameter $a$, and the concavity parameter $b$, as shown in figure 8 (details in section 5 of Supplemental Material).
@@
Figure 8: a) A set of bodies is defined by the parameters $a$ and $b$. b) As $a$ and $b$ defined are varied, we obtain settlers ($\alpha_{p}\alpha_{r}<0$), drifters ($\alpha_{p}\alpha_{r}=0$), and flutterers ($\alpha_{p}\alpha_{r}>0$). Increasing $b$ turns settlers into flutterers, but increasing $a$ at fixed $b$ eventually turns flutterers back into settlers.
We use the Boundary Integral Method [5, 25, 26, 27, 28] to obtain $\cal R$ numerically. The parameters $\alpha_{p},\,\alpha_{r},\,G_{1},\,G_{2},\,G_{3}$ completely describe the sedimentation problem in a quiescent fluid, with figure 8 showing their dependence on shape. For a given $b$, an optimal $a$ maximizes $\alpha_{p}\alpha_{r}$, resulting in the highest fluttering frequency.
To summarize, we have shown that di-bilaterals show three kinds of Stokesian sedimentation, depending only on their shape, appearing through the parameter $\alpha_{p}\alpha_{r}$. Settlers and drifters asymptotically fall vertically and obliquely respectively, without rotating. In flutterers, periodic dynamics in $\bm{p}_{y}$ drives a periodic or quasi-periodic response of $\bm{p}_{x}$, based on the ratio $\mu/\omega_{y}$. Both time-scales depend only on the conserved quantity $H$. All three classes of bodies can be obtained by bending a capsule about its central axis and making its cross-section concavo-convex. Previous studies on asymmetric bodies have well-understood $\bm{p}_{y}$ dynamics, predicting fixed points of differing character in [23, 16, 17, 6]. Among these, we rule out limit cycles for di-bilaterals by additional conserved quantity, $H$. Except for a qualitative argument which argues for of non-chaotic dynamics [13], the $\bm{p}_{x}$ dynamics has not been studied before to our knowledge. The $\bm{p}_{x}$ dynamics enables an understanding and quantitative description of quasi-periodicity. We also extend the proof [16] of no persistent drift of torque-free skewed bodies to any arbitrary body which performs periodic/quasi-periodic motion.
Acknowledgements.
@@
We thank Saumav Kapoor for help in designing flutterers. We acknowledge support of the Department of Atomic Energy, Government of India, under project no. RTI4001. We thank the referees for their suggestions, which improved the paper.
@@ Appendix B Discrete and continuous symmetries of the rotational dynamics of flutterers
The discrete symmetry group of the rotational dynamical equations (4) and (8) in the main paper is $\mathcal{G}_{R}=\mathbb{Z}_{2}\times\mathbb{Z}_{2}\times\mathbb{Z}_{2}$. The group $\mathcal{G}_{R}$ is an abelian group containing 8 elements which corresponds to the following eight transformations of dynamical variables:
@@
The conserved quantity $H$ is obtained as a first integral using: $\bm{Q}_{1}\cdot\bm{\Lambda}_{2}=\alpha_{p}+\alpha_{r}+t\dfrac{d}{dt}\log(p_{1y}^{\alpha_{p}}p_{3y}^{\alpha_{r}})$. The other scalar products of $\bm{Q}$’s and $\bm{\Lambda}$’s yield the constraints (9) in the main paper as the conserved quantities.
Appendix C Details of dynamics of flutterers
C.1 Persistent horizontal drift
In this section we ask if a di-bilateral, which is a flutterer, can ever show persistent drift in the horizontal plane and answer the question separately for the quasiperiodic and the periodic cases. For the former, following Thorp & Lister [15] we exploit the discrete symmetries of the rotational dynamical equations to rule out persistent drift.
@@ C.1 Persistent horizontal drift
Claim: Flutterers can have persistent drift in the plane perpendicular to gravity only if the driving frequency $\omega_{y}$ of the $\bm{p}_{y}$ dynamics is exactly equal to the response frequency $\mu$ of the $\bm{p}_{x}$ dynamics, i.e., $\mu=\omega_{y}$. In this case, the Poincare map $\mathcal{P}_{t_{0}}^{t_{0}+T_{y}}$ is just the identity map.
@@
Proof: The rotational dynamical equations (4) and (8) in the main paper have a discrete symmetry group $\mathcal{G}_{R}=\mathbb{Z}_{2}\times\mathbb{Z}_{2}\times\mathbb{Z}_{2}$. If a given rotational trajectory is invariant under $g\in\mathcal{G}_{R}$ and $g(v_{i})=-v_{i}$ then there can be no persistent drift along the $x_{i}$ direction for that trajectory, for the drift at one point is exactly cancelled by the other mapped point. Note that $g:(p_{1y},p_{2y},p_{3y},p_{1x},p_{2x},p_{3x})\to(p_{1y},p_{2y},p_{3y},-p_{1x},-p_{2x},-p_{3x})$ is a symmetry of equations (4) and (8) in the main paper with $g(v_{x})=-v_{x}$ and $g(v_{z})=-v_{z}$. Thus the invariance of trajectories of the rotational dynamical system under $g$ is sufficient to rule out any persistent horizontal drift.
@@
For quasi-periodic cases, any point $\bm{p}_{x}(t)$ in a given trajectory will be mapped to the point $-\bm{p}_{x}$ at time $t+kT_{y}$, for some $k\in\mathbb{N}$. This is due to the fact that the Poincare map $\mathcal{P}_{t}^{t+T_{y}}$ fills up the entire great circle with normal vector $\bm{p}_{y}(t)$. Since this great circle has an inversion symmetry, there exist a $k\in\mathbb{N}$ such that $\bm{p}_{x}(t+kT_{y})=-\bm{p}_{x}$. Therefore, the region occupied by quasi-periodic orbits on $S^{2}$ has inversion symmetry $\bm{p}_{x}\to-\bm{p}_{x}$. This rules out any persistent drift in the horizontal plane owing to the discrete symmetry element $g\in\mathcal{G}_{R}$.
@@
For the periodic case we have closed trajectories subject to constraints (9) in the main paper which may not have the inversion symmetry $\bm{p}_{x}\to-\bm{p}_{x}$ (see figure (6) in the main paper). The discrete symmetry argument to rule out persistent drift is inconclusive in this case. A necessary condition for persistent drift can be obtained as follows: owing to equation (11) in the main paper, each term on the right hand side of $v_{x}$ and $v_{z}$ in equation (12) in the main paper is of the form $e^{i\mu t}f(t)$, where $f(t)=f(t+T_{y})$ is some $T_{y}$ periodic function. The contribution of such terms to the displacement over one time period $T$ of a periodic trajectory is
@@
In section I.2 we show how persistent drift can be ruled out for arbitrary bodies which performs periodic/quasi-periodic motion, except for the special case where the time period of $\bm{p}_{y}$ dynamics exactly matches that of the $\bm{p}_{x}$ dynamics.
For flutterers, numerics suggests that the ratio $\mu/\omega_{y}$ is bounded between the interval [11/2, $[1-1/\sqrt{2},\,0.5]$ in which case there is no persistent drift. The lower limit of the ratio can be obtained by linearizing $\bm{p}_{y}$ dynamics near $H=H_{\text{max}}$, in which case the time period of $\bm{p}_{y}$ dynamics is periodic with time period $T_{ym}=2\pi/\sqrt{2\alpha_{p}\alpha_{r}}$. The $\bm{p}_{x}$ dynamics upto the zeroth order in $H-H_{\text{max}}$ is given by $\bm{\dot{p}}_{x}=\bm{A}^{*}\bm{p}_{x}$, where $\bm{A}^{*}$ is the matrix in equation (8) of the main paper, evaluated at one of the centre fixed points of the $\bm{p}_{y}$ dynamics. Therefore, upto the zeroth order the $\bm{p}_{x}$ dynamics is also periodic with the time period $T_{xm}=2\pi/\sqrt{\alpha_{p}\alpha_{r}}$. Now, using equation (11) of the main paper, $\bm{p}_{x}(t+T_{ym})=\bm{p}_{x}(t)e^{i\mu T_{ym}}$; and since $\bm{p}_{x}(t)\sim e^{\pm 2\pi it/T_{xm}}$, we have $e^{\pm 2\pi iT_{ym}/T_{xm}}=e^{i\mu T_{ym}}$, which gives $\mu T_{ym}/2\pi=1-1/\sqrt{2}$.
When $H=H_{\text{max}}$, we still have no persistent drift, since the valid solutions of $\bm{p}_{x}:\bm{p}_{x}(t)=e^{\bm{A}^{*}t}\bm{p}_{x}(0)$ are periodic with zero mean. This is because the eigenvector of $\bm{A}^{*}$ with zero eigenvalue does not satisfy the constraint $\bm{p}_{x}\cdot\bm{p}_{y}=0$ and thus cannot be part of the solution.
Finally, for $H=0$, the body exponentially reaches one of the saddle points $\bm{p}_{y}^{*}=(0,\pm 1,0)$ along its stable manifold as $t\to\infty$ and consequently the horizontal drift goes to zero exponentially.
C.2 Chirality of translational trajectories
We discuss the chirality of the translational trajectories of flutterers. As the body settles, its motion in the horizontal $x-z$ plane is either clockwise or anti-clockwise when viewed from below, corresponding to negative or positive chirality respectively. In figure 3 of the main paper, regions on $\mathcal{S}^{2}$ with $p_{1y}p_{3y}>0$ correspond to different chirality in translation motion from regions with $p_{1y}p_{3y}<0$. This can be seen as follows: for any given angular trajectory ($\bm{p}_{y},\bm{p}_{x}$) with $p_{1y}p_{3y}>0$, the transformation $g:(p_{1y},\,p_{1x})\to(-p_{1y},\,-p_{1x})$ maps the original trajectory to the trajectory in the region $p_{1y}p_{3y}<0$ with the same $H$. Thus the mapped trajectory has same time period $T_{y}$ in $\bm{p}_{y}$ dynamics and same frequency ratio $\mu/\omega_{y}$ in $\bm{p}_{x}$ dynamics. However, $g:v_{z}\to-v_{z}$, while the other velocity components remain invariant. So the trajectory of the translational dynamics ($x(t),y(t),z(t)$) is mapped to the trajectory $(x(t),y(t),-z(t))$, i.e., the original and mapped trajectories are of opposite chirality.